How many kg of water must be evaporated from 50kg of a 10% salt solution to obtain a 15% salt solution?
This was given to me out of a GEPA math tutorial book.

Answer by stanbon(75887)

How many kg of water must be evaporated from 50kg of a 10% salt solution to obtain a 15% salt solution?
-----------
Amt of water in 10% solution = 0.90*50kg = 45kg
---------
Let amt of water evaporated be "x" kg.
--------------
Amt of water in 15% solution = 0.85(50-x)=(42.5-0.85x)kg
--------------
EQUATION:
water - water = water
45 - x= 42.5-0.85x
2.5=0.15x
x= 6 2/3 kg (amt of water that must be evaporated)
==================
Cheers,
Stan H.